# Risks of Practicing Large Language Models in Smart Grid: Threat Modeling and Validation

### v2

@@ Abstract
Large language models (LLMs) represent significant breakthroughs in artificial intelligence and hold considerable potential for applications within smart grids. However, as demonstrated in previous literature, AI technologies are susceptible to various types of attacks. It is crucial to investigate and evaluate the risks associated with LLMs before deploying them in critical infrastructure like smart grids. In this paper, we systematically evaluated the risks of LLMs and identified two major types of attacks relevant to potential smart grid LLM applications, presenting the corresponding threat models. We also validated these attacks using popular LLMs and real smart grid data. Our validation demonstrates that attackers are capable of injecting bad data and retrieving domain knowledge from LLMs employed in different smart grid applications.
@@ I Introduction
Artificial Intelligence (AI) technologies have been actively researched for different applications in smart grids. In recent years, various AI techniques ranging from classical Support Vector Machines (SVM) [1] to Artificial Neural Networks (ANN) [2] have been proposed for diverse applications within smart grids, including anomaly detection [2], load forecasting [3], energy theft detection [4], and so on.
Recently, a significant breakthrough in AI has been the development of Large Language Models (LLMs). LLMs have a large number of model parameters and possess the capability to process general natural language [5]. Due to their exceptional performance across various natural language processing (NLP) tasks, LLMs have facilitated numerous applications across different industries [6, 7]. Given the successful performance of AI techniques in numerous smart grid applications, it is natural to explore the potential of integrating LLMs into the smart grid [8, 9, 10, 11].
@@
Despite the outstanding performance of AI models, they can also be vulnerable and expose systems to new types of cyberattacks. Previous research has shown that popular AI models, such as deep neural networks, are susceptible to various attacks, including adversarial [12] and backdoor attacks [13]. These attacks have proven to be effective against AI applications in smart grids as well [14, 15, 16, 17]. Therefore, conducting thorough research on potential risks is crucial before deploying LLMs in smart grid applications.
The applications of LLMs are rapidly evolving, and research into their risks, especially in potential smart grid applications, is still nascent. To the best of our knowledge, [10] represents the sole research focusing on the security of LLMs in smart grid contexts; it outlines potential threat types but lacks a comprehensive threat model and validation. To address this gap, we systematically study the potential threats posed by LLMs in smart grid applications, develop generalized threat models, and validate these threats through simulated LLM attacks. Our contributions are summarized as follows:
  • We analyze how the threats posed by LLMs differ from those associated with previous AI models (e.g. SVM) and identify the scope of risks that LLMs bring to the smart grid.
  • Considering attack motivations, we propose two general threat models for potential attacks on LLMs deployed in smart grids: (1) bad data injection attacks by external adversaries, and (2) domain knowledge extraction attacks by internal actors.
  • As a proof of concept, we validate these two threat types using popular LLMs: OpenAI GPT-3.5 [18], OpenAI GPT-4 [19], and Meta LLaMA-3 [20]. Through thorough simulations, we demonstrate that attackers can readily inject malicious data into, and extract confidential information from, these models.
@@ II-A Machine Learning and NLP in Power Grids
Machine learning techniques have been extensively studied and applied in various smart grid applications, including load forecasting [22], anomaly detection [23], energy theft detection [4], and cyberattack detection [2, 24, 25, 26].
In recent years, natural language processing (NLP) applications have also appeared in smart grids. In 2016, Sun et al. utilized social network information to detect power outages in specific areas [27]. They employed Bayesian models to analyze Twitter messages, achieving accurate detection results. Similarly, Wang et al. [23] collected public news and utilized a pre-trained language model to detect incidents at renewable energy facilities.
In addition, research into the application of LLMs to power systems has begun. Dong et al. [11] explored potential use cases of LLMs within power systems, demonstrating the capabilities of fine-tuned LLMs in pattern recognition tasks such as equipment damage detection and document analysis. Similarly, Huang et al. [9] examined the use of LLMs in specific power system operations, including optimal power flow (OPF) and electric vehicle (EV) scheduling. These studies highlight the potential utility of LLMs in enhancing the efficiency and reliability of smart grids.
Further research on the application of machine learning in power grids can be found in [28].
II-B Machine Learning Security in Smart Grid
While ML has achieved outstanding results in many smart grid tasks, its security remains a significant concern.
Well-performing ML models have been shown to be vulnerable to adversarial attacks, a phenomenon originally demonstrated in the field of computer vision [12]. This type of attack has also proven effective against ML-based smart grid applications. In 2018, Chen et al. first explored adversarial examples with both categorical and sequential applications in power systems [14]. Later, Li et al. [17] demonstrated that attackers could craft adversarial data capable of bypassing ML-based cyberattack detection while adhering to the physical constraints of the power grid. [29] showed that adversarial attackers could also circumvent ML-based energy theft detection models while reporting extremely low power consumption data.
@@ III-B Threat Analysis: Scope
  • General LLM Inherent Limitations: Attacks such as jailbreaks, which attempt to circumvent LLM restrictions to output sensitive content like personal information, represent general issues for all LLM applications and are not unique to smart grids.
@@ III-C Threat Models
Although not explicitly mentioned, most AI methods proposed for smart grid applications in previous literature assume that the AI model is trained by trusted personnel, such as internal engineers, and that all data used for training comes from reliable sources like smart meter data [22, 23, 4, 2]. However, this assumption no longer holds when applying LLMs. Due to the extremely high data requirements and costs associated with LLM training, pre-trained models are typically used. These pre-trained LLMs, whether open-source (e.g., Meta LLaMA-3) or proprietary (e.g., OpenAI GPT-4), have pre-trained parameters that may lead to unexpected behavior in response to well-crafted inputs from attackers.
@@ III-C2 Domain Knowledge Extraction
  • Attacker Capabilities: An insider attacker can interact with the LLM, including sending inputs and receiving outputs.
@@
We focus on insider attackers in the context of domain knowledge extraction, as best practices recommend avoiding the provision of sensitive information to publicly accessible LLMs. Thus, insider attackers, who have legitimate access, can interact with the LLM by sending inputs and receiving outputs. By analyzing these outputs, the attacker can derive detailed information about the domain knowledge, potentially extracting sensitive information not intended for disclosure.
@@ IV Validation
In this section, we validate the two types of attacks—bad data injection and domain knowledge extraction—by simulating scenarios involving LLM applications in a smart grid context. We selected the proprietary models OpenAI GPT-3.5 and GPT-4, as well as the open-source model Meta LLaMA-3, for our study due to their popularity and extensive coverage in related research [11, 9, 37]. Our validation results demonstrate that, under the practical threat models proposed in this paper, attackers can successfully execute both types of attacks on potential smart grid LLM applications.
The evaluations of the OpenAI models were implemented through the official OpenAI API [38]. For Meta LLaMA-3, we deployed the LLaMA-3.1-8B-Instruct [39] model on Amazon Web Services (AWS) using Hugging Face Inference Endpoints. Due to page limitations, only essential results are presented here; the implementation details, such as the complete system prompts and the datasets used in this paper, can be found in [21].
@@ IV-A2 Attack Simulation
  • Phase 1 - Normal Inputs: We reproduce the binary classification task using the LLMs. We empirically drafted a system prompt to identify relevant messages, and the same prompt was used for all three models. A sample of the system prompt can be found in Fig. 4. In this phase, we queried all three LLMs with normal messages, and the final input to the LLMs can be represented as
@@
    Similar to Phase 1, we queried the LLMs with the same messages but interspersed them with the three types of bad data sequentially to observe the effects on LLM responses. The performances of the LLMs under bad data injection attacks are shown in Table. II. By comparing the performances of the LLMs on normal inputs and injection inputs, we observe a significant drop in associated metrics for all three LLMs under all three different attack scenarios. These evaluation results demonstrate that well-designed injection texts can deceive the LLMs into misclassifying the inputs, highlighting the vulnerability of LLMs to bad data injection attacks in potential smart grid applications.
@@ IV-B1 Application Introduction
To demonstrate the vulnerability of LLMs to domain knowledge extraction attacks from insider attackers, we consider the use case where LLMs are employed by a utility company acting as a virtual assistant for internal employees. Specifically, the LLMs are prompted with raw smart grid data, which we consider as domain knowledge, but are intended to respond to employees with only aggregated information about the data. Under normal circumstances, the LLMs would reject requests to output detailed data.
@@ IV-B2 Attack Simulation
  • Phase 1 Normal Query: We directly challenge the LLMs by requesting them to output all the data they possess. This phase involves 100 tests for all three LLMs, each employing different groups of data (domain knowledge) to assess the robustness of the LLMs’ data confidentiality mechanisms under straightforward query conditions.
    In our evaluation, all three models successfully adhered to their configuration by not providing detailed data across all 100 tests. This demonstrates their effectiveness in safeguarding sensitive information under straightforward querying conditions. A sample of normal query input and LLM response is presented in Fig. 6
  • Phase 2 Domain Knowledge Extraction: In the attack phase, we employ well-crafted queries designed to subtly coax the LLMs into revealing more information than they are configured to disclose. This phase also consists of 100 tests using the same groupings of data as in Phase 1 for all three LLMs, to measure the LLMs’ resilience against more sophisticated extraction techniques.
    When faced with well-crafted, nuanced inputs designed to elicit more information than they are configured to disclose, all three models failed to maintain their data prevention protocols in all tests. In these cases, all models disclosed all the provided data, highlighting a vulnerability in their ability to handle sophisticated domain knowledge extraction attacks. The sampling output of the LLMs is shown in Fig. 6.

### v3

@@ Abstract
Large language models (LLMs) represent significant breakthroughs in artificial intelligence and hold potential for applications within smart grids. However, as demonstrated in previous literature, AI technologies are susceptible to various types of attacks. It is crucial to investigate and evaluate the risks associated with LLMs before deploying them in critical infrastructure like smart grids. In this paper, we systematically evaluated the risks of LLMs and identified two major types of attacks relevant to potential smart grid LLM applications, presenting the corresponding threat models. We validated these attacks using popular LLMs and real smart grid data. Our validation demonstrates that attackers are capable of injecting bad data and retrieving domain knowledge from LLMs employed in different smart grid applications.
@@ I Introduction
Artificial Intelligence (AI) technologies have been extensively studied for a range of smart grid applications, including anomaly detection [1], load forecasting [2], and energy theft detection [3]. A recent milestone in AI is the development of Large Language Models (LLMs). Unlike traditional task-specific AI applications—which require data collection, model design and training, and inference—LLMs are pretrained on general knowledge and can be directly applied to downstream tasks through prompts. This significantly simplifies and accelerates the development of AI applications [4]. Due to their remarkable performance, LLMs have enabled a wide range of applications across various industries [5, 6]. Given the proven success of AI techniques in smart grid contexts, it is natural to explore the potential integration of LLMs into smart grid systems [7, 8, 9, 10, 11].
@@
Despite the outstanding performance of AI models, they can also be vulnerable and expose systems to cyberattacks. Previous research has shown that popular AI models, such as Deep Neural Networks (DNNs), are susceptible to various attacks like adversarial attacks [12]. These attacks have proven to be effective against AI applications in smart grids as well [13, 14, 15, 16]. Therefore, conducting thorough research on potential risks is crucial before deploying LLMs in smart grid applications.
The applications of LLMs are rapidly evolving, and research into their risks, especially in potential smart grid applications, is still nascent. [10] studies the security of LLMs in smart grid contexts; it outlines potential threat types but lacks a comprehensive threat model and validation. To address this gap, we systematically study the potential threats posed by LLMs in smart grid applications, develop generalized threat models, and validate these threats through simulated LLM attacks. Our contributions are summarized as follows:
  • We analyze how the threats posed by LLMs differ from those associated with previous AI models and identify the scope of risks that LLMs bring to the smart grid.
  • Considering attack motivations, we propose two general threat models for potential attacks on LLMs deployed in smart grids: (1) bad data injection attacks by external adversaries, and (2) domain knowledge extraction attacks by internal & external actors.
  • As a proof of concept, we validate these two threat types using popular LLMs: OpenAI GPT-3.5 [17], OpenAI GPT-4 [18], Meta LLaMA-3 [19], Google Gemini 2.0 [20], and DeepSeek-V3 [21]. Through thorough simulations, we demonstrate that attackers can readily inject malicious data into, and extract confidential information from all these LLMs.
@@ II-A Machine Learning and NLP in Power Grids
Machine learning (ML) techniques have been extensively studied and applied in various smart grid applications, including load forecasting [23], anomaly detection [24], energy theft detection [3], and cyberattack detection [1, 25]. In recent years, Natural Language Processing (NLP) applications have also appeared in smart grids. In 2016, Sun et al. utilized social network information to detect power outages in specific areas [26]. They employed Bayesian models to analyze Twitter messages, achieving accurate detection results. Similarly, Wang et al. [24] collected public news and utilized a pre-trained language model to detect incidents at renewable energy facilities.
Research on the application of LLMs in power systems has emerged. Dong et al. [11] investigated potential use cases of LLMs in power systems, demonstrating the effectiveness of fine-tuned LLMs in pattern recognition tasks such as equipment damage detection and document analysis. Similarly, Huang et al. [9] explored the use of LLMs in specific power system operations, including optimal power flow (OPF) and electric vehicle (EV) scheduling. In [27], LLMs were applied to power dispatching tasks and demonstrated superior performance compared to existing strategies. These studies highlight the promising role of LLMs in improving the efficiency and reliability of smart grids.
II-B AI Security in Smart Grid
While ML has achieved outstanding results in many smart grid tasks, its security remains a significant concern [13, 16, 28, 29]. The well performed ML models have been shown to be vulnerable to adversarial attacks. In 2018, Chen et al. first explored adversarial examples with both categorical and sequential applications in power systems [13]. Later, [16] demonstrated that attackers could craft adversarial data capable of bypassing ML-based cyberattack detection while adhering to the physical constraints of the power grid. [29] showed that adversarial attackers could also circumvent ML-based energy theft detection models while reporting extremely low power consumption data.
@@ III-B Threat Analysis: Scope
  • General LLM Inherent Limitations: Attacks such as jailbreaks [34], which attempt to circumvent LLM restrictions to output sensitive content like personal information, represent general issues for all LLM applications and are not unique to smart grids.
@@ III-C Threat Models
Although not explicitly mentioned, most AI methods proposed for smart grid applications in previous literature assume that the AI model is trained by trusted personnel, such as internal engineers, and that all data used for training comes from reliable sources like smart meter data [23, 24, 3, 1]. However, this assumption no longer holds when applying LLMs. Due to the extremely high data requirements and costs associated with LLM training [37], pre-trained models are typically used. These pre-trained LLMs, whether open-source (e.g., Meta LLaMA-3) or proprietary (e.g., OpenAI GPT-4), have pre-trained parameters that may lead to unexpected behavior in response to well-crafted inputs from attackers.
@@ III-C2 Domain Knowledge Extraction
  • Attacker Capabilities: An attacker can interact with the LLM, including sending inputs and receiving outputs.
@@
In practice, insider attackers are more common under the context of domain knowledge extraction, as best practices recommend avoiding the provision of sensitive information to publicly accessible LLMs. However, the threat model can also be extended to outsider attackers under certain assumptions—for example, if a user-facing LLM chatbot is misconfigured to access unauthorized data.
The attacker who have legitimate access, can interact with the LLM by sending inputs and receiving outputs. By analyzing these outputs, the attacker can derive detailed information about the domain knowledge, potentially extracting sensitive information not intended for disclosure.
@@ IV Validation
In this section, we validate the two types of attacks—bad data injection and domain knowledge extraction—by simulating scenarios involving LLM applications in a smart grid context. We selected the proprietary models OpenAI GPT-3.5, GPT-4 and Google Gemini 2.0, as well as the open-source model Meta LLaMA-3 and DeepSeek-V3, for our study due to their popularity and extensive coverage in related research [11, 9, 39]. Our validation results demonstrate that, under the practical threat models proposed in this paper, attackers can successfully execute both types of attacks on potential smart grid LLM applications.
The evaluations of the OpenAI, Google, and DeepSeek models were implemented through the associated official APIs. For Meta LLaMA-3, we deployed the LLaMA-3.1-8B-Instruct [40] model on Amazon Web Services (AWS) using Hugging Face Inference Endpoints. Due to page limitations, only essential results are presented here; the implementation details, such as the complete system prompts and the datasets used in this paper, can be found in [22].
@@ IV-A2 Attack Simulation
  • Phase 1 - Normal Inputs: We reproduce the binary classification task using the LLMs. We empirically drafted a system prompt to identify relevant messages, and the same prompt was used for all five models. A sample of the system prompt can be found in Fig. 4. In this phase, we queried all LLMs with normal messages, and the final input to the LLMs can be represented as
@@
    Similar to Phase 1, we queried the LLMs with the same messages but interspersed them with the three types of bad data sequentially to observe the effects on LLM responses. The performances of the LLMs under bad data injection attacks are shown in Table. II. By comparing the performances of the LLMs on normal inputs and injection inputs, we observe a significant drop in associated metrics for all five LLMs under all three different attack scenarios. These evaluation results demonstrate that well-designed injection texts can deceive the LLMs into misclassifying the inputs, highlighting the vulnerability of LLMs to bad data injection attacks in potential smart grid applications.
@@
Google Gemini 2.0 Accuracy 59.3% 52.2% 34.0%
Precision 54.0% 0% 35.2%
Recall 100% 0% 45.5%
F1 70.1% 0% 39.7%
DeepSeek-V3 Accuracy 48.0% 52.2% 35.0%
Precision 47.9% 0% 40.1%
Recall 100% 0% 72.8%
F1 64.7% 0% 51.7%
@@ IV-B1 Application Introduction
To demonstrate the vulnerability of domain knowledge extraction attacks, we consider the use case where LLMs are employed within the control center as a virtual assistant. Specifically, the LLMs are prompted with raw smart grid data, which we consider as domain knowledge, but are intended to respond to users with only aggregated information about the data. Under normal circumstances, the LLMs would reject requests to output detailed data.
@@ IV-B2 Attack Simulation
  • Phase 1 Normal Query: We directly challenge the LLMs by requesting them to output all the data they possess. This phase involves 100 tests for all five LLMs, each employing different groups of data (domain knowledge) to assess the robustness of the LLMs’ data confidentiality mechanisms under straightforward query conditions.
    In our evaluation, all models successfully adhered to their configuration by not providing detailed data across all 100 tests. This demonstrates their effectiveness in safeguarding sensitive information under straightforward querying conditions. A sample of normal query input and LLM response is presented in Fig. 6
  • Phase 2 Domain Knowledge Extraction: In the attack phase, we employ well-crafted queries designed to subtly coax the LLMs into revealing more information than they are configured to disclose. This phase also consists of 100 tests using the same groupings of data as in Phase 1 for all LLMs, to measure the LLMs’ resilience against more sophisticated extraction techniques.
    When faced with well-crafted, nuanced inputs designed to elicit more information than they are configured to disclose, all models failed to maintain their data prevention protocols in all tests. In these cases, all models disclosed all the provided data, highlighting a vulnerability in their ability to handle sophisticated domain knowledge extraction attacks. The sampling output of the LLMs is shown in Fig. 6.